A critical reaction in the production of energy to do work or drive chemical reactions in biological systems is the hydrolysis of adenosine triphosphate, ATP, to adenosine diphosphate, ADP, as described by

ATP(aq) + H2O(l) --> ADP(aq) + HPO4(aq)

or which ΔG°rxn = –30.5 kJ/mol at 37.0 °C and pH 7.0. Calculate the value of ΔGrxn in a biological cell in which [ATP] = 5.0 mM, [ADP] = 0.50 mM, and [HPO42–] = 5.0 mM.

Delta Grxn =
Answer #1

Given:-

standard free energy change (\DeltaG0rxn) = - 30.5 KJ/mol = - 30500 J /mol

free energy change (\DeltaGrxn) = ?

As we know that

biological temperature (T) = 37 0C = 273 + 37 = 310 K

Gas constant (R) = 8.314 JK-1mol-1

molar concentration ATP i.e [ATP] = 5.0 mM = 5.0 \times 10-3 M

molar concentration ADP i.e [ADP] = 5.0 mM = 5.0 \times 10-3 M

molar concentration HPO42- i.e [ HPO42-] = 5.0 mM = 5.0 \times 10-3 M

Since we know that

ATP(aq) + H2O(l)  \rightleftharpoons ADP(aq) + HPO42-(aq)

therefore

Equilibrium constant (Keq) = [ADP][ HPO42-] / [ATP][ H2O]

Equilibrium constant (Keq) = 5.0 \times​​​​​​​ 10-3\times 5.0 \times​​​​​​​ 10-3  / 5.0 \times​​​​​​​ 10-3\times 1 (since [ H2O] = 1)

Equilibrium constant (Keq) = 25.0 \times​​​​​​​ 10-6   / 5.0 \times​​​​​​​ 10-3

Equilibrium constant (Keq) = 5.0 \times​​​​​​​ 10-3

As we know that according to the formula

free energy change (\DeltaGrxn) =  (\DeltaG0rxn) + 2.303RT \times (Keq)

free energy change (\DeltaGrxn) =   - 30500 J /mol   + ( 2.303 \times 8.314 JK-1mol-1\times 310 K \times 5.0 \times​​​​​​​ 10-3 )

free energy change (\DeltaGrxn) =   - 30500 J /mol   + ( 29678.07 \times​​​​​​​ 10-3 J / mol )

free energy change (\DeltaGrxn) =   - 30500 J /mol   + ( 29.67807 J / mol )

free energy change (\DeltaGrxn) =   - 30500 J /mol   + 29.67807 J / mol

free energy change (\DeltaGrxn) =   - 30470.32 J /mol  

free energy change (\DeltaGrxn) =   - 30.47032 KJ /mol (i.e the answer)